# Algorithm Design and Prototype Validation for Reconfigurable Intelligent Sensing Surface: Forward-Only Transmission

### v1

@@ IV-A Hardware Introductions
This section introduces the hardware employed within the system and delineates its implementation throughout. Fig. 3 presents an overview of the system’s hardware, while Table-II elaborates on the detailed parameters and functions of these hardware components. Note that we utilize an additional external clock (i.e., GPSDO) to synchronize local frequencies across different USRPs.
@@ IV-A3 Reflecting hardware
Reconfigurable Intelligent Surface (RIS) 430 mm (W) $\times$ 430 mm (H) $\times$ 46 mm (L) 2-bits phase adjustable. $10\times 10$ elements. Frequency band $3.4\sim 3.6$ GHz.
Sensing Antenna Array 236 mm(W) $\times$ 59 mm (H) $\times$ 9.2 mm (L) Four elements linear array microstrip antenna. Frequency band $3.08\sim 3.91$ GHz. Total gain 12.6 dBi. Element spacing $0.6883\lambda$, where $\lambda$ is the wavelength.
@@ IV-C Sensing Model
We emphasize again that our primary aim is to selectively receive the desire signal from the mixed signal through the assistance of sensing. The sensing model comprises two tasks: DOA estimation and the identification recognition of the two transmitters.
@@ IV-C1 DOA Estimation
$\mathbf{x}(n)=\mathbf{A}\mathbf{s}(n)+\mathbf{v}(n),$ (29)
where $\mathbf{x}(n)\in\mathbf{C}^{N_{a}\times 1}$ is the received signals from sensing array at snapshot $n$, $\mathbf{A}=\left[\bm{\alpha}\left(\phi_{\text{Tx2R}}^{1}\right),\bm{\alpha}\left(\phi_{\text{Tx2R}}^{2}\right)\right]\in\mathbb{C}^{N_{a}\times 2}$ is the direction matrix, and $\bm{\alpha}\left(\phi_{\text{Tx2R}}^{i}\right)=\left[1,e^{-\mathbbm{i}\phi_{\text{Tx2R}}^{i}},\cdots,e^{-\mathbbm{i}(N_{a}-1)\phi_{\text{Tx2R}}^{i}}\right]^{T},\forall i\in\{1,2\}$ is the direction vectors. $\mathbf{s}(n)\in\mathbb{C}^{2\times 1}$ and $\mathbf{v}(n)\in\mathbb{C}^{N_{a}\times 1}$ denote the transmit signals and zero-mean additive white Gaussian noise (AWGN) vector at snapshot $n$ with the normalized noise power $\sigma_{0}^{2}$.
@@ V Experiment Results
In this section, we practical verify the performance of RISS-assisted system. The entire system including two single horn-antenna Tx (i.e., $M=1$) and one single horn-antenna Rx, which are both controlled by the USRP X300 and transmit/receive single-carrier QPSK signals, as demonstrated in Section IV-A to IV-B. The location of the center of the RISS is set to $[0,0,0]$. The distance between Tx/Rx and the center of the RISS is set to $d_{c}=3.5$ meters, and the Tx/Rx and RISS are on the same plane, which means that $\phi^{\text{azi}}=\pi/2$, thus, the location of the Tx and Rx is set to $\mathbf{b}_{\text{Tx},i}=[d_{c}\sin(\phi^{\text{ele}}_{i}),0,d_{c}\cos(\phi^{\text{ele}}_{i})],\forall i\in\{1,2\}$ and $\mathbf{b}_{\text{Rx}}=[0,0,d_{c}]$. In the experiments, we can obtain $\phi^{\text{ele}}_{i}$ from the DOA algorithm, e.g., MUSIC and ROOTMUSIC, and the identification can be also obtained and determine which Tx is the target source, as depicts in Section IV-C.
@@ V-A Sensing Performance
Figure 8: The experiment results of both MUSIC and ROOTMUSIC implemented by the sensing antenna array within the RISS. The Tx. 1 is located in $\phi^{\text{ele}}_{1}=-20^{\circ}$ and Tx. 2 is located in $\phi^{\text{ele}}_{2}=10^{\circ}$.
@@
    Spatial filtering: This process is used to separate signals from the combined received data, resulting in distinct samples for the target and interference signals.
Following the pre-processing steps, two signal sets are obtained and used for binary classification with SR-net. The SR-net is implemented using PyTorch [37] and trained on a single NVIDIA RTX A6000 GPU, utilizing the Adam optimizer [38] for 100 epochs. After training the model, it is deployed offline on the upper computer connected to the RISS for sensing.
@@ V-B Sensing-assisted Communication for Signal Enhancement
In this section, we first shows the sensing-assisted communication with only one Tx, i.e., $\tau=+\infty$ in Eq. (27) and $\eta=0$ in Eq. (28) for simplification. This means that our sole objective in this section is to enhance the target signal without considering interference suppression. Fig. 9 demonstrates that RISS enhance the communication performance although the phase shifts of the RISS are not appropriate, since the RISS can increase multipath and change the polarization direction of the signal [^7] The RISS reverses the polarization, resulting in the polarization of the Tx and Rx being opposite..
(a) RISS off.
(b) RISS on.
@@
Fig. 10 demonstrates the constellation diagram measurements of the receiver. Note that perfect sensing assumes knowledge of the direction of Tx. 2, i.e., $\phi_{2}^{\text{ele}}=10^{\circ}$. In contrast, imperfect sensing refers to cases where the angle is derived from actual sensing results. As observed in Fig. 10a and 10c, both near-field and far-field models enhance communication performance and significantly improve the SINR at the receiver compared to Fig. 9b. This improvement enables the adoption of higher-order modulation schemes, such as 16QAM and 64QAM, which enhance communication efficiency.
@@
(a) Near-field model with perfect sensing.
(b) Near-field model with imperfect sensing.
(c) Far-field model with perfect sensing.
(d) Far-field model with imperfect sensing.
@@
Figure 11: The normalized beampattern of near-field model and far-field model. The objective of both scheme is to maximize the signal power at the Tx with the knowledge of $\phi_{2}^{\text{ele}}=10^{\circ}$.
@@ V-C Sensing-assisted Communication for Interference Suppression
In this section, we further evaluate the performance of interference suppression. The directions of the target signal and interference signal are set to $\phi^{\text{ele}}_{1}=-20^{\circ}$ and $\phi^{\text{ele}}_{2}=10^{\circ}$, respectively. As shown in Fig. 11, the interference signal is located in the side-lobe direction, leading to strong interference at the receiver. Fig. 12 presents the communication constellation diagram with interference suppression using the near-field model, compared to the counterpart scheme that only aligns with the target signal. It can be observed that the interference signal reduces the SINR and causes the constellation diagram to become more dispersed. Furthermore, the constellation diagram exhibits a pattern similar to that of non-orthogonal multiple access (NOMA), due to the power disparity between Tx. 1 and Tx. 2. However, the interference suppression scheme significantly improves the SINR, restoring the compactness of the constellation diagram. The beampatterns of both schemes are shown in Fig. 13. It is important to note that the interference suppression scheme slightly reduces the main-lobe gain while significantly decreasing the side-lobe gain associated with the interference.
(a) Only align the target signal.
(b) Robust interference suppression for interference signal.
@@
Figure 13: The beampattern of align scheme and interference suppression scheme in the near-field model. The target signal is located in $\phi_{1}^{\text{ele}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{ele}}=10^{\circ}$.
In contrast, as illustrated in Fig. 14 and Fig. 15, the interference suppression in the far-field model is weak. This phenomenon may be attributed to the relatively close distance, rendering the far-field model inapplicable, or possibly due to RIS phase errors leading to interference suppression failure.
(a) Only align the target signal.
(b) Robust interference suppression for interference signal.
@@
Figure 15: The beampattern of align scheme and interference suppression scheme in the far-field model. The target signal is located in $\phi_{1}^{\text{ele}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{ele}}=10^{\circ}$.
Finally, to further investigate interference suppression, we deactivate Tx. 2, allowing the receiver to capture only the interference signal. The constellation measurements for both the align scheme and the interference suppression scheme using the near-field model are displayed in 16. With only Tx. 1 active, the constellation in 16a shows that the align scheme retains significant interference signal power. In contrast, the interference suppression scheme shows effective interference mitigation, resulting in a highly dispersed constellation diagram of the interference signal, further demonstrating the robustness of the interference suppression technique.
(a) Interference constellation with align scheme.
(b) Interference constellation with robust interference suppression.

### v2

@@ IV-A Hardware Introductions
This section introduces the hardware employed within the system and delineates its implementation throughout. Fig. 4 presents an overview of the system’s hardware, while Table-II elaborates on the detailed parameters and functions of these hardware components. Note that we utilize an additional external clock (i.e., GPSDO) to synchronize local frequencies across different USRPs. Additional hardware parameters and descriptions, including the detailed specifications of the 2-bit RIS, the three-dimensional structural model, the $S_{11}$ reflection coefficient curve, and the radiation pattern of the sensing antenna array, are available in [39], which may offer further insights.
@@ IV-A3 Reflecting hardware
Reconfigurable Intelligent Surface (RIS) 430 mm (W) $\times$ 430 mm (H) $\times$ 46 mm (L) 2-bits phase adjustable. $10\times 10$ elements. Frequency band $3.4\sim 3.6$ GHz. Element spacing 42.83 mm, i.e., $0.5\lambda$, where $\lambda$ is the wavelength at 3.5GHz.
Sensing Antenna Array 236 mm(W) $\times$ 59 mm (H) $\times$ 9.2 mm (L) Four elements linear array microstrip antenna. Frequency band $3.08\sim 3.91$ GHz. Total gain 12.6 dBi. Element spacing 58.96 mm, i.e., $0.6883\lambda$, where $\lambda$ is the wavelength at 3.5GHz.
@@ IV-C Sensing Model
We emphasize again that our primary aim is to selectively receive the target signal from the mixed signal through the assistance of sensing. The sensing model comprises two tasks: DOA estimation and the identification recognition of the two transmitters.
@@ IV-C1 DOA Estimation
$\mathbf{x}(n)=\mathbf{A}\mathbf{s}(n)+\mathbf{n}_{z}(n),$ (30)
where $\mathbf{x}(n)\in\mathbf{C}^{N_{a}\times 1}$ is the received signals from sensing array at snapshot $n$, $\mathbf{A}=\left[\bm{\alpha}\left(\phi_{\text{Tx2R}}^{1}\right),\bm{\alpha}\left(\phi_{\text{Tx2R}}^{2}\right)\right]\in\mathbb{C}^{N_{a}\times 2}$ is the direction matrix, and $\bm{\alpha}\left(\phi_{\text{Tx2R}}^{i}\right)=\left[1,e^{-\mathbbm{i}\phi_{\text{Tx2R}}^{i}},\cdots,e^{-\mathbbm{i}(N_{a}-1)\phi_{\text{Tx2R}}^{i}}\right]^{T},\forall i\in\{1,2\}$ is the direction vectors. $\mathbf{s}(n)\in\mathbb{C}^{2\times 1}$ and $\mathbf{n}_{z}(n)\in\mathbb{C}^{N_{a}\times 1}$ denote the transmit signals and zero-mean additive white Gaussian noise (AWGN) vector at snapshot $n$ with the normalized noise power $\sigma_{0}^{2}$.
@@ V Experiment Results
In this section, we practical verify the performance of RISS-assisted system. The entire system including two single horn-antenna Tx (i.e., $M=1$) and one single horn-antenna Rx, which are both controlled by the USRP X300 and transmit/receive carrier QPSK signals, as demonstrated in Section IV-A to IV-B. The location of the center of the RISS is set to $[0,0,0]$. The distance between Tx/Rx and the center of the RISS is set to $d_{c}=3.5$ meters, and the Tx/Rx and RISS are on the same plane, which means that $\phi^{\text{ele}}=\pi/2$, thus, the location of the Tx and Rx is set to $\mathbf{b}_{\text{Tx},i}=[d_{c}\sin(\phi^{\text{azi}}_{i}),0,d_{c}\cos(\phi^{\text{azi}}_{i})],\forall i\in\{1,2\}$ and $\mathbf{b}_{\text{Rx}}=[0,0,d_{c}]$. In the experiments, we can obtain $\phi^{\text{azi}}_{i}$ from the DOA algorithm, e.g., MUSIC and ROOTMUSIC, and the identification can be also obtained and determine which Tx is the target source, as depicts in Section IV-C. The interference suppression factors are set to $\tau=0.1$ (i.e., $-10$ dB) and $\eta=1$, respectively [^10] Note that excessively large values of $\tau$ and $\eta$ may lead to a greater emphasis on interference suppression, potentially at the expense of target signal enhancement. Therefore, when selecting the values of both parameters, it is essential to strike a balance between interference suppression and the enhancement of the target signal, in accordance with the specific requirements of the system.. Moreover, the Error Vector Magnitude (EVM) is adopted as a key metric to assess the modulation accuracy in digital communication systems. Typically expressed as a percentage, EVM quantifies the deviation between the received signal $s_{received}$ and the ideal reference signal $s_{ideal}$. It is mathematically defined as
$\Delta_{\text{EVM}}(\%)=\frac{RMS(s_{received}-s_{ideal})}{RMS(s_{ideal})}\times 100\%,$ (39)
where RMS()($\cdot$)( ⋅ ) denotes the root mean square operation. The RMS values for the error vector and the ideal signal are given by
$RMS(s_{received}-s_{ideal})=\sqrt{\frac{1}{d}\sum_{d=1}^{|\mathcal{D}|/2}\left|s_{received,d}-s_{ideal,d}\right|^{2}},$
$RMS(s_{ideal})=\sqrt{\frac{1}{d}\sum_{d=1}^{|\mathcal{D}|/2}\left|s_{ideal,d}\right|^{2}},$ (40)
where $|\mathcal{D}|/2$ denotes the number of modulated symbols, given that the $|\mathcal{D}|$ bits are modulated using QPSK with a modulation order of 2.
@@ V-A Sensing Performance
Figure 9: The experiment results of both MUSIC and ROOTMUSIC implemented by the sensing antenna array within the RISS. The Tx. 1 is located in $\phi^{\text{azi}}_{1}=-20^{\circ}$ and Tx. 2 is located in $\phi^{\text{azi}}_{2}=10^{\circ}$.
@@
    Spatial filtering: This process is used to separate signals from the combined received data, resulting in distinct samples for the target and interference signals. Specifically, the samples are weighted and summed individually to implement spatial filtering based on the DOA results, thereby distinguishing and obtaining two separate sets of signal samples from the spatial characteristics. For further details, please refer to the MVDR [34] and Linearly Constrained Minimum Variance (LCMV)[43] methods.
Following the pre-processing steps, two signal sets are obtained and used for binary classification with SR-net. The SR-net is implemented using PyTorch [44] and trained on a single NVIDIA RTX A6000 GPU, utilizing the Adam optimizer [45] for 100 epochs. After training the model, it is deployed offline on the upper computer connected to the RISS for sensing. The detialed processes are depicted in Fig. 10.
Figure 10: The processes of DOA estimation, spatial filtering and RF fingerprint recognition.
@@ V-B Sensing-assisted Communication for Signal Enhancement
In this section, we first shows the sensing-assisted communication with only one Tx, i.e., $\tau=+\infty$ in Eq. (28) and $\eta=0$ in Eq. (29) for simplification. This means that our sole objective in this section is to enhance the target signal without considering interference suppression [^11] For detailed implementation videos, please refer to https://drive.google.com/file/d/11PmI_luWSJ0AxLBtrOmpnz8PKYNMoFP4/view?usp=drive_link.. Fig. 11 demonstrates that RISS enhance the communication performance although the phase shifts of the RISS are not appropriate, since the RISS can increase multipath and change the polarization direction of the signal [^12] The RISS reverses the polarization, resulting in the polarization of the Tx and Rx being opposite..
(a) RISS off, $\Delta_{\text{EVM}}=44.7494\%$.
(b) RISS on, $\Delta_{\text{EVM}}=26.4598\%$.
@@
Fig. 12 demonstrates the constellation diagram measurements of the receiver. Note that perfect sensing assumes knowledge of the direction of Tx. 2, i.e., $\phi_{2}^{\text{azi}}=10^{\circ}$. In contrast, imperfect sensing refers to cases where the angle is derived from actual sensing results. As observed in Fig. 12a and 12c, both near-field and far-field models enhance communication performance and significantly improve the SINR at the receiver compared to Fig. 11b. This improvement enables the adoption of higher-order modulation schemes, such as 16QAM and 64QAM, which enhance communication efficiency.
@@
(a) Near-field model with perfect sensing, $\Delta_{\text{EVM}}=4.1357\%$.
(b) Near-field model with imperfect sensing, $\Delta_{\text{EVM}}=4.6768\%$.
(c) Far-field model with perfect sensing, $\Delta_{\text{EVM}}=5.1486\%$.
(d) Far-field model with imperfect sensing, $\Delta_{\text{EVM}}=5.2401\%$.
@@
Figure 13: The normalized beampattern of near-field model and far-field model. The objective of both scheme is to maximize the signal power at the Tx with the knowledge of $\phi_{2}^{\text{azi}}=10^{\circ}$.
@@ V-C Sensing-assisted Communication for Interference Suppression
In this section, we further evaluate the performance of interference suppression [^13] For detailed implementation videos, please refer to https://drive.google.com/file/d/11SJT2nxz9Zuo5t7ZtBZR4MCcfASvaUpO/view?usp=drive_link.. The directions of the target signal and interference signal are set to $\phi^{\text{azi}}_{1}=-20^{\circ}$ and $\phi^{\text{azi}}_{2}=10^{\circ}$, respectively. As shown in Fig. 13, the interference signal is located in the side-lobe direction, leading to strong interference at the receiver. Fig. 14 presents the communication constellation diagram with interference suppression using the near-field model, compared to the counterpart scheme that only aligns with the target signal. It can be observed that the interference signal reduces the SINR and causes the constellation diagram to become more dispersed. Furthermore, the constellation diagram exhibits a pattern similar to that of non-orthogonal multiple access (NOMA), due to the power disparity between Tx. 1 and Tx. 2. However, the interference suppression scheme significantly improves the SINR, restoring the compactness of the constellation diagram. The theoretical and experimental beampatterns of both schemes are shown in Fig. 15. It is noteworthy that both the theoretical and experimental results presented in Fig. 15a and Fig. 15b demonstrate a consistent trend in interference suppression and target signal enhancement achieved through the 2-bit RISS. Specifically, a gain is observed in the direction of the target signal, while deep nulls are formed in the directions of interference. However, due to practical system imperfections such as phase shift deviations in the reflection elements and environmental disturbances, the actual level of interference suppression exhibits a noticeable deviation from the theoretical prediction. Additionally, it is also observed that the interference suppression scheme results in a slight reduction in the main-lobe gain compared to the only align scheme.
(a) Only align the target signal, $\Delta_{\text{EVM}}=19.2229\%$.
(b) Robust interference suppression for interference signal, $\Delta_{\text{EVM}}=6.5897\%$.
@@
(a) Theoretical beam gain with 2-bit RISS.
(b) Practical experimental beam gain.
Figure 15: The beampattern of align scheme and interference suppression scheme in the near-field model. The target signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$.
In contrast, as illustrated in Fig. 16 and Fig. 17, the performance of target enhancement and interference suppression in the far-field model is weak. To better understand the underlying cause of this phenomenon, we present theoretical 2D heatmaps under both the NF and FF schemes. It can be observed from Fig. 18 that both schemes place the Tx.2 within the main-lobe while forming nulls in the directions of Tx.1. However, the NF scheme exhibits fewer sidelobes and does not generate significant signal gain in the near region. In contrast, the FF scheme produces more pronounced sidelobes and exhibits gain along the entire path at $\phi_{2}^{azi}=10^{\circ}$. In addition, there is a slight offset between the main-lobe direction and the actual position of the Tx.2. This misalignment is caused by the combined effects of the interference suppression constraint and the 2-bit RISS quantization. Furthermore, hardware imperfections, such as phase adjustment errors in the RISS, may also be contributing factors to the inferior performance of the FF scheme compared to the NF scheme.
(a) Only align the target signal, $\Delta_{\text{EVM}}=43.5608\%$.
(b) Robust interference suppression for interference signal, $\Delta_{\text{EVM}}=20.0157\%$.
@@
(a) Theoretical beam gain with 2-bit RISS.
(b) Practical experimental beam gain.
Figure 17: The beampattern of align scheme and interference suppression scheme in the far-field model. The target signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$.
(a) Theoretical heatmap of NF scheme.
(b) Theoretical heatmap of FF scheme.
Figure 18: Heatmaps of both NF and FF schemes. The RISS is located at (0,0) and target signal is located in $\phi_{1}^{\text{azi}}=-20^{\circ}$, while the interference signal is located in $\phi_{2}^{\text{azi}}=10^{\circ}$.
Finally, to further investigate interference suppression, we deactivate Tx. 2, allowing the receiver to capture only the interference signal. The constellation measurements for both the align scheme and the interference suppression scheme using the near-field model are displayed in 19. With only Tx. 1 active, the constellation in 19a shows that the align scheme retains significant interference signal power. In contrast, the interference suppression scheme shows effective interference mitigation, resulting in a highly dispersed constellation diagram of the interference signal, further demonstrating the robustness of the interference suppression technique.
(a) Interference constellation with align scheme, $\Delta_{\text{EVM}}=24.5761\%$.
(b) Interference constellation with robust interference suppression, $\Delta_{\text{EVM}}=59.7931\%$.
@@ VI Conclusion
Appendix A Hardware Information Supplement
The detailed parameters of RIS can be found in Table V.
TABLE V: Product specifications for the 3.5 GHz RIS from Beijing Actenna Technology Co., Ltd.
Product Specifications
Product Model I24-S35
Working Frequency 3.4–3.6 GHz
Array Size 10$\times$10 units
Phase-Shifting Step 2-bit
Phase Accuracy 90$\pm$15
Insertion Loss 1.0–3.0 dB
Polarization Mode Orthogonal dual-linear polarization
Angle Range $\pm$60 (H/V)
Power Consumption 9W (Typical)
Control Port DB9
Input Voltage 9–36V DC
Dimensions 430 mm (W) $\times$ 430 mm (H) $\times$ 46 mm (D)
Weight 3.0 kg
Moreover, Fig. 20 illustrate the three-dimensional structural model of the sensing antenna array, which consists of four rectangular microstrip patch antennas arranged in parallel, with a back-fed configuration. To mitigate antenna coupling and address other design considerations, the spacing between the units is set to 0.0590 meters, corresponding to 0.6883 wavelengths at 3.5 GHz. The physical dimensions of the entire antenna array are 236 mm × 59 mm × 9.2 mm.
Figure 20: Three-dimensional structural model of the sensing antenna array.
Figure 21: $S_{11}$ reflection coefficient curve of sensing antenna array.
Fig. 21 shows the $S_{11}$ reflection coefficient curve. Using a reference of -10 dB, which corresponds to a reflection loss of less than or equal to 10%, the antenna array exhibits a bandwidth ranging from approximately 3.09 GHz to 3.91 GHz, resulting in a total bandwidth of about 830 MHz. The resonance point is observed around 3.58 GHz, suggesting that the antenna array is capable of operating effectively at 3.5 GHz, as discussed in this paper.
Figure 22: The beampattern of sensing antenna array.
Fig. 22 presents the radiation pattern of the sensing antenna array at an operating frequency of 3.5 GHz. The measured maximum gain reaches 12.6 dB, demonstrating strong directional radiation capability within the main lobe. As illustrated in Fig. 20, the $y$-axis is defined as the horizontal axis in this configuration, with horizontal linear polarization aligned along the $y$-axis direction.